what are the equations for each of the conic sections?
ex. circle is (y-h)^2+(x-k)^2=radius^2

Answer by venugopalramana(3286)

PARABOLA
(X-H)^2=4A(Y-K)......WHERE AXIS IS X=H , VERTEX IS (H,K),FOCUS IS (H,K+A),ECCENTRICITY IS 1 AND DIRECTRIX IS Y=K-A
OR..(Y-K)^2=4A(X-H)..WHERE AXIS IS Y=K , VERTEX IS (H,K),FOCUS IS (H+A,K)ECCENTRICITY IS 1, AND DIRECTRIX IS X=H-A
ELLIPSE
{(X-H)^2/A^2}+{(Y-K)^2/B^2}=1..WHERE AXIS IS X=H , CENTRE IS (H,K),ECCENTRICITY IS <1

HYPERBOLA
{(X-H)^2/A^2}-{(Y-K)^2/B^2}=1...WHERE AXIS IS X=H , CENTRE IS (H,K),ECCENTRICITY IS >1